# A polynomial delay algorithm generating all potential maximal cliques in triconnected planar graphs

### v2

@@ 4 Polynomial delay generation of PMCs of a triconnected planar graph
A natural approach to generate $\Pi^{\prime}(G)$ is to generate all minimal separators and, for each $S$ output with $|S|\geq 4$, generate PMCs that contain $S$ based on the proof of Theorem 3.16, which shows that $X\supset S$ is a PMC if and only if $L_{G}[X]$ is an $(S,P)$-steering, where $P=X\setminus S$. To implement this approach, we need to deal with the problem of suppressing duplicate outputs of a single element without introducing super-polynomial delay. We use the following technique due to Bergougnoux, Kanté and Wasa [1] to address this problem.
@@
Under these assumptions, Algorithm 1 generates ${\mathcal{S}}(G)$ with polynomial delay. The idea of this algorithm is as follows. For each $s\in{\mathcal{S}}(G)$, let us say that $i\in I$ owns $s$ if $i$ is the largest $j$ such that $s\in{\mathcal{S}}_{j}(G)$. We let ${\textsc{Gen}}_{i}$, where $i$ owns $s$, be responsible for the output of $s$ and suppress the output of $s$ from ${\textsc{Gen}}_{j}$ for other $j$. The executions of ${\textsc{Gen}}_{i}$ for $i\in I$ are scheduled in such a way that at most a polynomial number of outputs are suppressed before an unsuppressed output occurs.
@@
Algorithm 1 generates ${\mathcal{S}}(G)$ with polynomial delay, if all of the four conditions above are satisfied.
Proof 4.6.
@@
Definition 4.7.
Let $G$ be a triconnected plane graph and $C$ a minimally separated component of $G$ with $|S|\geq 4$ where $S=N_{G}(C)$. Recall that there are exactly two full components associated with $S$ and let $C^{\prime}$ be the full component distinct from $C$. A port of $S$ in $C^{\prime}$ is a vertex $u\in C^{\prime}$ such that $N_{L_{G}}(u)\cap S$ is a slot of $L_{G}[S]$. For a socket $\{R_{1},R_{2}\}$ of $L_{G}[S]$, a port pair on $\{R_{1},R_{2}\}$ in $C^{\prime}$ is an unordered pair $\{u_{1},u_{2}\}$ of ports of $S$ in $C^{\prime}$ such that $N_{L_{G}}(u_{i})\cap S=R_{i}$ for $i\in\{1,2\}$.
Lemma 4.8.
Let $G$ be a triconnected plane graph and $C$ a minimally separated component of $G$ with $|S|\geq 4$, where $S=N_{G}(C)$. Let $C^{\prime}$ be the other full component of $S$. Let $P$ be an arbitrary subset of $C^{\prime}$ with $|P|\geq 2$ and let $X=S\cup P$. Then, $L_{G}[X]$ is an $(S,P)$-steering if and only if there is a socket $\{R_{1},R_{2}\}$ of $L_{G}[S]$ and a port pair $\{u_{1},u_{2}\}$ in $C^{\prime}$ on this socket such that either $L_{G}[P]$ is a chordless path of $A(C,u_{1},u_{2})$ between $u_{1}$ and $u_{2}$, where $A(C,u_{1},u_{2})$ is the subgraph of $L_{G}$ induced by the vertex set $(C^{\prime}\setminus N_{L_{G}}(S))\cup\{u_{1},u_{2}\}$, or there is a hinge $s$ of the socket $\{R_{1},R_{2}\}$ and $L_{G}[P]$ is a chordless path of $B(C,u_{1},u_{2},s)$ between $u_{1}$ and $u_{2}$, where $B(C,u_{1},u_{2},s)$ is the subgraph of $L_{G}$ induced by the vertex set $(C^{\prime}\setminus N_{L_{G}}(S\setminus\{s\}))\cup\{u_{1},u_{2}\}$.
Proof 4.9.
Suppose $L_{G}[X]$ is an $(S,P)$-steering. Let $H=L_{G}[X]$. Due to Definition 3.1, $H[P]=L_{G}[P]$ is a path. Let $u_{1}$ and $u_{2}$ be the two ends of this path and let $R_{i}=N_{L_{G}}(u_{i})\cap S$ for $i\in\{1,2\}$. Since $\{R_{1},R_{2}\}$ is a socket of the cycle $L_{G}[S]$, $R_{i}$ is a slot of $L_{G}[S]$ for $i\in\{1,2\}$ and, moreover, $R_{1}\cup R_{2}$ is not a slot of $L_{G}[S]$. Furthermore, $N_{L_{G}}(P\setminus\{u_{1},u_{2}\})\cap S$ is either empty or a singleton $\{s\}$ where $s$ is a hinge of the socket $\{R_{1},R_{2}\}$. Suppose the former is the case. Then, $L_{G}[P]$ is a path in $A(C,u_{1},u_{2})$ since no internal vertex of it is adjacent in $L_{G}$ to any vertex in $S$. It is a chordless path, since if it had a chord in $A(C,u_{1},u_{2})$, then the graph $L_{G}[P]$ would contain that chord and $L_{G}[P]$ would not be a path. Now suppose the latter, that $N_{L_{G}}(P\setminus\{u_{1},u_{2}\})\cap S$ is a singleton $\{s\}$ where $s$ is a hinge of the socket $\{R_{1},R_{2}\}$. Then, $L_{G}[P]$ is a path in $B(C,u_{1},u_{2},s)$ since no internal vertex of it is adjacent in $L_{G}$ to any vertex in $S\setminus\{s\}$.
For the other direction, suppose that there is a socket $\{R_{1},R_{2}\}$ of $L_{G}[S]$ and a port pair $\{u_{1},u_{2}\}$ on this socket such that either (1) $L_{G}[P]$ is a chordless path of $A(C,u_{1},u_{2})$ between $u_{1}$ and $u_{2}$ or (2) there is a hinge $s$ of the socket $\{R_{1},R_{2}\}$ and $L_{G}[P]$ is a chordless path of $B(C,u_{1},u_{2},s)$ between $u_{1}$ and $u_{2}$. If (1) is the case, then $N_{L_{G}}(P\setminus\{u_{1},u_{2}\})\cap S$ is empty. If, on the other hand, (2) is the case, then $N_{L_{G}}(P\setminus\{u_{1},u_{2}\})\cap S$ is either empty or a singleton $\{s\}$. In either case, $L_{G}[S\cup P]$ is an $(S,P)$-steering.
For a minimally separated component $C$ of a triconnected plane graph $G$ with $|S|\geq 4$ where $S=N_{G}(C)$, let $\Pi(G,C)$ denote the set of PMCs $X$ such that $C\in{\mathcal{C}}_{G}(X)$.
@@
Let $G$ and $C$ be as above. Then, $\Pi(G,C)\neq\emptyset$.
Proof 4.11.
@@
Proof 4.13.
Due to the proof of Theorem 3.16, $X\subseteq V(G)$ belongs to $\Pi(G,C)$ if and only if $L_{G}[X]$ is an $(S,P)$-steering where $P=X\setminus S$. Based on Lemma 4.8, $\Pi(G,C)$ can be generated as follows. First assume that $|S|\geq 5$. Let $C^{\prime}$ be the other full component of $S=N_{G}(C)$. For every socket $\{R_{1},R_{2}\}$ of $L_{G}[S]$ and every port pair $\{u_{1},u_{2}\}$ on $\{R_{1},R_{2}\}$ in $C^{\prime}$, do the following. If the socket does not have a hinge, then generate all chordless paths of $A(C,u_{1},u_{2})$ between $u_{1}$ and $u_{2}$, where $A(C,u_{1},u_{2})$ is as defined in Lemma 4.8; if the socket has a hinge $s$ then generate all chordless paths of $B(C,u_{1},u_{2},s)$ between $u_{1}$ and $u_{2}$, where $B(C,u_{1},u_{2},s)$ is as defined in Lemma 4.8. For each path $p$ generated, output $S\cup V(p)$. Note that, in the latter case, the hinge $s$ of the socket is unique since $|S|\geq 5$ and hence PMCs in $\Pi(G,C)$ are generated without duplication.
If $|S|=4$, a socket may have two hinges. We apply the technique of Theorem 4.14 to deal with this situation. Let $S=\{s_{1},s_{2},s_{3},s_{4}\}$ and suppose these vertices appear on $L_{G}[S]$ in this clockwise order. For each $i\in\{1,2,3,4\}$, let $\{R_{i,1},R_{i,2}\}$ denote the socket of $L_{G}[S]$ of which $s_{i}$ is a hinge. Observe that such a socket is unique for fixed $i$. Let $\Pi_{i}(G,C)$ be the set of PMCs $X\in\Pi(G,C)$ such that $L_{G}[X\setminus S]$ is a path of $L_{G}$ between $u_{1}$ and $u_{2}$, where $\{u_{1},u_{2}\}$ is a port pair on the socket $\{R_{i,1},R_{i,2}\}$. It is straightforwardly confirmed below that the four subsets $\Pi_{i}(G,C)$, $i\in\{1,2,3,4\}$, of $\Pi(G,C)$ satisfy the four conditions for Algorithm 1 to be applicable. Let $n=V(G)$.
@@
    For each $X\in\Pi(G,C)$ and $i\in\{1,2,3,4\}$, it can be decided whether $X\in\Pi_{i}(G,C)$ in time $n^{O(1)}$.
@@
    For each $i\in\{1,2,3,4\}$, there is an algorithm that generates $\Pi_{i}(G,C)$ with polynomial delay.
Therefore, Algorithm 1 applied to this situation generates $\Pi(G,C)$ with polynomial delay.
@@
Proof 4.15.
It suffices to show that $\Pi^{\prime}(G)$, the set of PMCs $X$ of $G$ such that $|N_{G}(C)|\geq 4$ for some $C\in{\mathcal{C}}_{G}(X)$, can be generated with polynomial delay. Let $n=|V(G)|$. For each $v\in G$, let $\Pi_{v}(G)$ denote the union of $\Pi(G,C)$ over all minimally separated components $C$ of $G$ such that $v\in C$ and $|N_{G}(C)|\geq 4$. We have $\Pi^{\prime}(G)=\bigcup_{v\in V(G)}\Pi_{v}(G)$ and $|V(G)|=n=n^{O(1)}$. Given $X\in\Pi(G)$, it is straightforward to test if $X\in\Pi_{v}(G)$ in polynomial time, for a fixed $v$. For each $v$, $\Pi_{v}(G)$ can be generated with polynomial delay: we generate all minimal separators $S$ such that $v\not\in S$ by the second part of Lemma 4.2 and, for each $S$ generated such that $|S|\geq 4$, generate $\Pi(G,C)$ for the full component $C$ of $S$ such that $v\in C$. The delay in this algorithm is polynomial, since minimal separators are generated with polynomial delay and, for each minimal separator generated, at least one PMC is generated due to Proposition 4.10.
Therefore, applying Lemma 4.5, we can generate $\Pi(G)$ with polynomial delay.
@@ 5 Computing the planar treewidth
An algorithm for generating PMCs immediately leads to a treewidth algorithm because of Theorem 2.2 due to Bouchitté and Todinca [7].
@@
Proof 5.2.
We first prove the case where $G$ is triconnected. We compute a planar embedding of $G$ in time $O(n)$. Then we apply our generation algorithm to compute $\Pi(G)$ in time $|\Pi(G)|n^{O(1)}$. Finally, we apply Theorem 2.2 to obtain ${\mathop{tw}}(G)$ in time $|\Pi(G)|n^{O(1)}$.
For a general planar graph $G$, let $s(G)$ denote the number of minimal separators of cardinality at most two. We prove the theorem by induction on $s(G)$. The base case $s(G)=0$, where $G$ is triconnected, is already proved. Suppose $s(G)>0$ and let $S$ be an arbitrary minimal separator of $G$ with $|S|\leq 2$. Let $C_{1}$, …, $C_{m}$ be the components of $G[V(G)\setminus S]$ and let $G_{i}=G[C_{i}\cup N_{G}(C_{i})]\cup K(N_{G}(C_{i}))$, for $i=1,\ldots,m$, be the graph obtained from $G[C_{i}\cup N_{G}(C_{i})]$, the subgraph of $G$ induced by the closed neighborhood of $C_{i}$, by filling the open neighborhood of $C_{i}$ into a clique. Then $S$ is a safe separator of $G$ in the sense of [5] and therefore we have ${\mathop{tw}}(G)=\max\{|S|,\max_{i\in\{1,\ldots,m\}}{\mathop{tw}}(G_{i})\}$. By the induction hypothesis, we can compute ${\mathop{tw}}(G_{i})$ for each $i$ in time $|\Pi(G_{i})|n^{O(1)}$. We have $\sum_{i}|\Pi(G_{i})|\leq|\Pi(G)|$, since every PMC of $G_{i}$ is a PMC of $G\cup K(S)$ and hence is a PMC of $G$ and, moreover, there is no PMC that is common to $G_{i}$ and $G_{j}$ for $i\neq j$. Therefore, we can compute ${\mathop{tw}}(G)$ in time $|\Pi(G)|n^{O(1)}$.

### v3

@@ 4 Polynomial delay generation of PMCs of a triconnected planar graph
To design an algorithm to generate all PMCs based on our characterization of PMCs stated in Theorem 3.15, we need some preparations.
Definition 4.5.
Let $G$ be a triconnected plane graph and $C$ a minimally separated component of $G$ with $|S|\geq 4$ where $S=N_{G}(C)$. Recall that there are exactly two full components associated with $S$ and let $C^{\prime}$ be the full component distinct from $C$. A port of $(S,C)$ is a vertex $u\in C^{\prime}$ such that $N_{L_{G}}(u)\cap S$ is a slot of $L_{G}[S]$. We call the slot $N_{L_{G}}(u)\cap S$ of $L_{G}[S]$ the slot for $u$. A pair of ports $u_{1}$ and $u_{2}$ is valid if the union of the slot for $u_{1}$ and the slot for $u_{2}$ is not a slot. A vertex $s\in S$ with two neighbors $s_{1}$ and $s_{2}$ on the cycle $L_{G}[S]$ is a hinge of a valid pair of ports $u_{1}$ and $u_{2}$ if the slot for $u_{1}$ is either $\{s_{1}\}$ or $\{s_{1},s\}$ and the slot for $u_{2}$ is either $\{s_{2}\}$ or $\{s_{2},s\}$. See Figure 5 for examples of ports, valid and invalid port pairs, and a hinge.
Figure 5: (A) White vertices are some ports of $(S,C)$ where $S$ consists of black vertices and $C$ is the full component associated with $S$ that lies in the outer face of the black cycle. (B) Some valid pairs of ports. The valid pair in the second example has a hinge $s$. (C) Some invalid pairs of ports.
Ports are potential ends of a path $p$ such that $L_{G}[X]$ for $X=S\cup V(p)$ is a steering. A pair of ports must be valid in order for the pair to be the ends of such path $p$. We generate such paths using the algorithm for generating chordless paths. The following lemma is the basis of this approach.
Lemma 4.6.
Let $G$ be a triconnected plane graph and $C$ a minimally separated component of $G$ with $|S|\geq 4$, where $S=N_{G}(C)$. Let $C^{\prime}$ be the other full component of $S$. Let $P$ be an arbitrary subset of $C^{\prime}$ with $|P|\geq 2$ and let $X=S\cup P$.
  1. 1.
    For each valid pair $u_{1}$ and $u_{2}$ of ports of $(S,C)$, let $A(C,u_{1},u_{2})$ be the subgraph of $L_{G}$ induced by the vertex set $(C^{\prime}\setminus N_{L_{G}}(S))\cup\{u_{1},u_{2}\}$. Then, $L_{G}[X]$ is an $(S,P)$-steering if and only if there is a valid pair $u_{1}$ and $u_{2}$ of the ports of $S$ such that $P=V(p)$ for some chordless path $p$ of $A(C,u_{1},u_{2})$ between $u_{1}$ and $u_{2}$.
  2. 2.
    For each valid pair $u_{1}$ and $u_{2}$ of ports of $(S,C)$ that has hinges and each hinge $s$ of the pair, let $B(C,u_{1},u_{2},s)$ be the subgraph of $L_{G}$ induced by the vertex set $(C^{\prime}\setminus N_{L_{G}}(S))\cup N_{L_{G}}(s)\cup\{u_{1},u_{2}\}$. Then, $L_{G}[X]$ is an $(S^{\prime},\{s\})$-steering for $s\in S$, where $S^{\prime}=X\setminus\{s\}$, if and only if there is a valid pair of ports $u_{1}$ and $u_{2}$ of $(S,C)$ with a hinge $s$ such that $P=V(p)$ for some chordless path $p$ of $B(C,u_{1},u_{2},s)$ between $u_{1}$ and $u_{2}$.
Proof 4.7.
For part 1, suppose $L_{G}[X]$ is an $(S,P)$-steering where $P\subseteq C^{\prime}$. Let $H=L_{G}[X]$. Due to Definition 3.1, $H[P]=L_{G}[P]$ is a path. Let $u_{1}$ and $u_{2}$ be the two ends of this path. Then, $u_{1}$ and $u_{2}$ are ports of $S$ and, moreover, the pair of these ports is valid. Since no internal vertex of the path $L_{G}[P]$ is adjacent in $L_{G}$ to any vertex in $S$, $L_{G}[P]$ is a path in $A(C,u_{1},u_{2})$. It is a chordless path, since if it had a chord in $A(C,u_{1},u_{2})$, then the graph $L_{G}[P]$ would contain that chord and $L_{G}[P]$ would not be a path.
For the other direction, suppose that there is a valid pair of ports $u_{1}$ and $u_{2}$ of $S$ such that $L_{G}[P]$ is a chordless path of $A(C,u_{1},u_{2})$ between $u_{1}$ and $u_{2}$. Then, $N_{L_{G}}(P\setminus\{u_{1},u_{2}\})\cap S$ is empty and hence $L_{G}[S\cup P]$ is an $(S,P)$-steering.
For part 2, suppose $L_{G}[X]$ is an $(S^{\prime},\{s\})$-steering where $s\in S$ and $S^{\prime}=X\setminus\{s\}$. Due to the proof of Lemma 3.11, $L_{G}[P]$ is a path of $L_{G}$ between some valid pair of ports $u_{1}$ and $u_{2}$ of $(S,C)$ and, moreover, $s$ is a hinge of this pair of ports. Since the only vertex in $S$ that can be adjacent to any internal vertex of $L_{G}[P]$ is $s$, $L_{G}[P]$ is a path in $B(C,u_{1},u_{2},s)$ between $u_{1}$ and $u_{2}$. It is a chordless path, since if it had a chord in $B(C,u_{1},u_{2},s)$, then the graph $L_{G}[P]$ would contain that chord and $L_{G}[P]$ would not be a path.
For the other direction, suppose that there is a valid pair of ports $u_{1}$ and $u_{2}$ of $S$ with a hinge $s$ such that $L_{G}[P]$ is a chordless path of $B(C,u_{1},u_{2},s)$ between $u_{1}$ and $u_{2}$. Then, $L_{G}[S^{\prime}]$, where $S^{\prime}=X\setminus\{s\}$, is a cycle and $L_{G}(X)$ is an $(S^{\prime},\{s\})$-steering.
In implementing the generation of PMCs based on this lemma, we need to deal with the problem of suppressing duplicate outputs of a single element without introducing super-polynomial delay. We use the technique due to Bergougnoux, Kanté and Wasa [1] to address this problem. They used this technique for a particular problem of generating what they call minimal disjunctive separators. We formulate their technique in the following theorem that can be used for wide range of generation problems.
@@
Theorem 4.8.
Under the above assumptions, ${\mathcal{S}}(G)$ can be generated with polynomial delay.
We prove this theorem by showing that Algorithm 1 generates ${\mathcal{S}}(G)$ with polynomial delay. The idea of this algorithm is as follows. For each $s\in{\mathcal{S}}(G)$, let us say that $i\in I$ owns $s$ if $i$ is the largest $j$ such that $s\in{\mathcal{S}}_{j}(G)$. We let ${\textsc{Gen}}_{i}$, where $i$ owns $s$, be responsible for the output of $s$ and suppress the output of $s$ from ${\textsc{Gen}}_{j}$ for other $j$. The executions of ${\textsc{Gen}}_{i}$ for $i\in I$ are scheduled in such a way that at most a polynomial number of outputs are suppressed before an unsuppressed output occurs.
@@
Algorithm˜1 generates ${\mathcal{S}}(G)$ with polynomial delay, if all of the four conditions above are satisfied.
Proof 4.10.
@@
For a minimally separated component $C$ of a triconnected plane graph $G$ with $|S|\geq 4$ where $S=N_{G}(C)$, let $\Pi(G,C)$ denote the set of PMCs $X$ such that $C\in{\mathcal{C}}_{G}(X)$ and $L_{G}[X]$ is a steering.
@@
Let $G$ be a triconnected plane graph and $C$ a minimally separated component of $G$ such that $|N_{G}(C)|\geq 4$. Then, $\Pi(G,C)\neq\emptyset$.
Proof 4.12.
@@
Proof 4.14.
Based on Lemma 4.6, we generate $\Pi(G,C)$ as follows. We first generate $X\in\Pi(G,C)$ such that $L_{G}[X]$ is an $(S,P)$-steering where $P=X\setminus S$. To do so, for every valid pair of ports $u_{1}$ and $u_{2}$ of $S$, we generate all chordless paths of $A(C,u_{1},u_{2})$ between $u_{1}$ and $u_{2}$, where $A(C,u_{1},u_{2})$ is as defined in Lemma 4.6, using the algorithm due to Sato and Uno given in Lemma 4.1. For each path $p$ generated, we output $S\cup V(p)$. Since the number of valid pairs of ports is $O(n^{2})$, the generation of $X$ in this category is with polynomial delay even if some valid pairs do not yield any such $X$.
We then generate $X\in\Pi(G,C)$ such that $L_{G}[X]$ is an $(S^{\prime},\{s\})$-steering where $s\in S$ and $S^{\prime}=X\setminus\{s\}$. To do so, for every valid pair of ports $u_{1}$ and $u_{2}$ of $S$ with hinges and for each hinge $s$ of this pair, we generate all chordless paths of $B(C,u_{1},u_{2},s)$ between $u_{1}$ and $u_{2}$. For each path $p$ generated, we output $S\cup V(p)$. The valid pair of $u_{1}$ and $u_{2}$ may have more than one ports when $|S|=4$, which could cause duplicate outputs of an identical $X$. We suppress such duplicate outputs using Theorem 4.8. Let $\mathcal{T}$} be the set of all triples $(u_{1},u_{2},s)$ where $(u_{1},u_{2})$ is a valid pair of ports of $(S,C)$ and $s$ is a hinge of this pair. Index the set $\mathcal{T}$} by $I=\{1,\ldots,N\}$ where $N=|{\mathcal{T}}|$. For each $i\in I$, let ${\mathcal{P}}_{i}$ denote the set of chordless paths of $B(C,u_{1},u_{2},s)$ between $u_{1}$ and $u_{2}$, where $(u_{1},u_{2},s)$ is the triple in $\mathcal{T}$} indexed by $i$. Let ${\mathcal{P}}=\bigcup_{i\in I}{\mathcal{P}}_{i}$. Our goal is to generate $\mathcal{P}$} without duplication. We verify the conditions for Theorem 4.8 to be applied.
@@
    For each path $p$ in $\mathcal{P}$} and each $i\in I$, it can be decided whether $p\in{\mathcal{P}}_{i}$ in time $n^{O(1)}$.
@@
    For each $i\in I$, we can generate ${\mathcal{P}}_{i}$ with polynomial delay as described above.
Therefore, Theorem 4.8 applies and we can generate $\mathcal{P}$} and hence all PMCs $X$ in this category, with polynomial delay.
Remark 4.15.
Since the PMCs $X$ in the second category in the above lemma, those such that $L_{G}[X]$ is an $(S^{\prime},\{s\})$-steering for some $s\in S$ where $S^{\prime}=X\setminus\{s\}$, also belong to $\Pi(G,C^{\prime})$ for some full component associated with $S^{\prime}$, it might appear that the cumbersome generation of those PMCs described in the above proof is unnecessary. However, if we generate only the members of $\Pi(G,C)$ in the first category, those $X$ such that $L_{G}[X]$ is an $(S,P)$-steering where $P=X\setminus S$, then it is possible for the generator of $\Pi(G,C)$ to produce no outputs. This is a problem, since there is no readily provable polynomial bound on the number of successive components $C$ for which the generation of $\Pi(G,C)$ produces no outputs.
@@
Proof 4.17.
It suffices to show that $\Pi^{\prime}(G)$, the set of PMCs $X$ of $G$ such that $|N_{G}(C)|\geq 4$ for some $C\in{\mathcal{C}}_{G}(X)$, can be generated with polynomial delay. Let $n=|V(G)|$. For each $v\in G$, let $\Pi_{v}(G)$ denote the union of $\Pi(G,C)$ over all minimally separated components $C$ of $G$ such that $v\in C$ and $|N_{G}(C)|\geq 4$. We have $\Pi^{\prime}(G)=\bigcup_{v\in V(G)}\Pi_{v}(G)$ and $|V(G)|=n=n^{O(1)}$. Given $X\in\Pi(G)$ and $v\in V(G)$, it is straightforward to test if $X\in\Pi_{v}(G)$ in polynomial time. For each $v$, $\Pi_{v}(G)$ can be generated with polynomial delay: we generate all minimal separators $S$ such that $v\not\in S$ by the second part of Lemma 4.2 and, for each $S$ generated such that $|S|\geq 4$, generate $\Pi(G,C)$ for the full component $C$ of $S$ such that $v\in C$. The delay in this algorithm is polynomial, since minimal separators are generated with polynomial delay and, for each minimal separator generated, at least one PMC is generated due to Proposition 4.11.
Therefore, applying Theorem 4.8, we can generate $\Pi(G)$ with polynomial delay.
@@ 5 Computing the planar treewidth
An algorithm for generating PMCs immediately leads to a treewidth algorithm for triconnected planar graphs because of Theorem 2.3 due to Bouchitté and Todinca [7]. To extend the algorithm for general planar graphs, we use the following result from [5]. A vertex set $S$ of a graph $G$ is an almost clique of $G$ if there is some $v\in S$ such that $S\setminus\{v\}$ is a clique of $G$.
Theorem 5.1 (Bodlaender and Koster [5]).
Let $G$ be a graph and $S$ a minimal separator of $G$ that is an almost clique. Let $C_{i}$, $1\leq i\leq m$, be the components of $G[V(G)\setminus S]$ and let $G_{i}=G[C_{i}\cup N_{G}(C_{i})]\cup K(N_{G}(C_{i}))$, for $i=1,\ldots,m$, be the graph obtained from $G[C_{i}\cup N_{G}(C_{i})]$, the subgraph of $G$ induced by the closed neighborhood of $C_{i}$, by filling the open neighborhood of $C_{i}$ into a clique. Then ${\mathop{tw}}(G)=\max\{|S|,\max_{i\in\{1,\ldots,m\}}{\mathop{tw}}(G_{i})\}$.
We use this theorem for the special case where $S$ is a two-vertex minimal separator: such $S$ is always an almost clique.
@@
Proof 5.3.
We assume that $G$ is biconnected. Otherwise, we compute the treewidth of each biconnected component and take the maximum.
We first prove the case where $G$ is triconnected. We compute a planar embedding of $G$ in time $O(n)$ [15]. Then we apply our generation algorithm to compute $\Pi(G)$ in time $|\Pi(G)|n^{O(1)}$. Finally, we apply Theorem 2.3 to obtain ${\mathop{tw}}(G)$ in time $|\Pi(G)|n^{O(1)}$.
For a general planar graph $G$, let $s(G)$ denote the number of two-vertex separators of $G$. We prove the theorem by induction on $s(G)$. The base case $s(G)=0$, where $G$ is triconnected, is already proved. Suppose $s(G)>0$ and let $S$ be an arbitrary two-vertex separator of $G$. Let $C_{1}$, …, $C_{m}$ be the components of $G[V(G)\setminus S]$. Since $G$ is biconnected, we have $N(C_{i})=S$ for every $i\in\{1,\dots,m\}$. Therefore, $S$ is a minimal separator of $G$ and, due to Theorem 5.1, we have ${\mathop{tw}}(G)=\max\{2,\max_{i\in\{1,\ldots,m\}}{\mathop{tw}}(G_{i})\}$ where $G_{i}=G[C_{i}\cup S]\cup K(S)$. Since $G_{i}$ is a subgraph of a minor of $G$ obtained by contracting $C_{i^{\prime}}\cup{v}$ into a single vertex for arbitrary $i^{\prime}\neq i$ and $v\in S$, $G_{i}$ is planar. Moreover, $G_{i}$ is biconnected since any cut vertex of $G_{i}$ would be a cut vertex of $G$. Therefore, we may apply the induction hypothesis to each $G_{i}$ and compute ${\mathop{tw}}(G_{i})$ for each $i$ in time $|\Pi(G_{i})|n^{O(1)}$. We claim that $\sum_{i}|\Pi(G_{i})|\leq|\Pi(G)|$. To see this, let $X$ be an arbitrary PMC of $G_{i}$ and $H$ a minimal triangulation of $G_{i}$ in which $X$ is a maximal clique. Let $H^{\prime}$ be an arbitrary minimal triangulation of $G$ such that $H^{\prime}[V(G_{i})]=H$. Since every minimal separator of $G$ that is a clique in a minimal triangulation of $G$ is a minimal separator of that triangulation [14], $S$ is a minimal separator of $H^{\prime}$. Therefore, $X$ is a maximal clique of $H^{\prime}$ and hence is a PMC of $G$. Moreover, $X$ cannot be a PMC of $G_{j}$ for any $j\neq i$ since $X$ is not a subset of $V(G_{j})$. Therefore, the claim holds. We conclude that we can compute ${\mathop{tw}}(G)$ in time $|\Pi(G)|n^{O(1)}$.
6 Future work
Our success in a special case of the open question whether $\Pi(G)$ can be computed in time $|\Pi(G)|n^{O(1)}$ does not seem to suggest any approach to solving the question for general graphs. Much more modest goal is to address the question for general planar graphs. It might be possible to extend our generation algorithm to general planar graphs by finding some way of handling PMCs that cross two-vertex minimal separators. Bart Jansen asked if the generation result could be extended to graphs embedded on tori. We observe that latching graphs are not appropriate tool for that purpose, since if $G$ is a triconnected graph embedded on a torus and $X$ is a PMC of $G$ then the subgraph of the latching graph of $G$, appropriately defined on tori, induced by $X$ may have edge crossings: the proof of Proposition 2.17 relies essentially on the assumption that $G$ is embedded in the sphere. Studying upper bounds on $|\Pi(G)|$ is another avenue of research. There is a class of planar graphs for which $|\Pi(G)|=\Omega(1.442^{n})$ [9]. These graphs are biconnected but not triconnected, so it would be interesting to ask if an upper bound smaller than this bound holds for triconnected planar graphs.